OpenStudy (anonymous):

x^2 + 1/x^2 = 3 Explain please.

OpenStudy (kinggeorge):

Multiply everything by \(x^2\) so you get \[x^4+1=3x^2\] Then set equal to 0 and factor \[x^4-3x^2+1=0 \longrightarrow (x^2-x-1) \times (x^2+x-1)\] From there, you can use the quadratic formula to solve.

OpenStudy (anonymous):

I think you factored wrong

OpenStudy (kinggeorge):

I don't think so...Wolfram seems to confirm me.

OpenStudy (anonymous):

You are correct. It must've been a different problem then, I cant remember the constant that came after the = sign, but I remember the polynomial that I got was prime, would that mean x is no solution?

OpenStudy (kinggeorge):

remind me what a prime polynomial is?